Can someone help me with this quetion?

An epidemiologic (al) investigation that was started on June 1, 2016 and lasted through May 31, 2018, identified a population of 1,700 individuals among whom 10 were found to have the disease under investigation. During the study period, 4 and 6 new cases of the disease were found on May 31, 2017 and November 30, 2017 respectively. 5 were reported deceased from other causes on January 31, 2018. 25 individuals were lost to follow-up by September 1, 2017. There were no additional deaths or losses to follow-up. Further analysis at the end of the investigative period also revealed that 13 of the diseased were exposed to asbestos, and 7 were not. Among the 1,680 non-diseased, 10 were exposed to asbestos, and others were not. Calculate the:

  1. Prevalence of disease in the period through May 30, 2017

  1. Prevalence of disease on November 30, 2017

  1. Cumulative incidence of disease at the end of the investigation
  2. Incidence density/incidence rate of disease at the end of the investigation
Answer #1

I. Prevalence rate:

Formula: total number of reported case/ total number of population*100

New case: 4

Old case: 10

Total: 14

= 16/1700*100

=0.82

Prevance rare= 0.82.

II. November

New case: 6

Old case: 10

Total : 16

Formula= 16/1700*100

=0.94

Prevalence= 0.94.

III. Cumulative incident:

Formula= total number of new case/ total population*100

New cases= 4+6+13= 23

= 23/1700*190

=1.35

Cumulative incidence= 1.35.

IV. Incident at end of study:

New case: 13 at the end of study.

So

= 13/1700*100

=0.76.